OpenStudy (anonymous):

Which statement is FALSE about the width of the central maximum of a circular-aperture diffraction pattern? [A]If the size of the aperture (D) is increased, then the circular spot increases in size.[B]If the distance to the screen (L) increases, then the circular spot increases in size.[C]The width of the central bright spot from a circular aperture is 1.22 times wider than the central bright band produced from a single-slit aperture if the circular aperture width has the same width as the single slit.[D]If the size of the aperture (D) is increased, then the circular spot decreases in size.

OpenStudy (anonymous):

The answer is either [A], [B], [C] or [D], it would not let me properly format the question so that's why it's in paragraph form. (Exceeded a word limit maybe?)

OpenStudy (anonymous):

Consider the equation for the size of an Airy disk:\[\sin \theta=\frac{ 1.22\lambda }{ d }\]where θ is the angular width of the Airy disk as measured from its center to the first zero (or minimum or fringe. It depends on what terminology you were taught to use); λ is the wavelength; and d is the aperture diameter. You can see mathematically that if d increases, the size (angular radius) of the spot decreases. That should help you find that answer.